A=1/2(a+b)h solve for a
Answer by stanbon(48545)

Multiply both sides by "2" to get:
2A=(a+b)h
Divide by "h" to get:
2A/h=a+b
Subtract "b" to get:
[2A/h]-b=a
Cheers,
stan H.